# Tightrope-walking rationality in action: Feyerabendian insights for the foundations of quantum mechanics

### v2

@@ 2 Relation between physics and philosophy, in the foundations of QM
[…] throughout the period of the older quantum theory Bohr underlined the ’formal’ character of the new ideas which arose in connection with the principle of correspondence: these ideas, we can read repeatedly in his earlier papers, lead to some surprising predictive successes, and they also establish some order in the steadily accumulating empirical material. However, they do not form a theory in the ordinary, or ’classical’ sense of the word, that is, they do not contain a coherent description of new and objective features. Nor can we say that they explain the relevant experimental facts. All this is emphasized by Bohr himself and not only once, but in almost every paper he writes. He also emphasizes that the dependence on classical ideas that is implied by the use of the correspondence principle characterizes ’the present state of science’, or ’the present state of our knowledge’, and not the essence of physics. It is even regarded as somewhat undesirable. In these early years Bohr’s aim seems to have been the gradual construction of an effective instrument of prediction which he hoped would act as a stimulus and as an aid towards the invention of a new atomic theory in the full sense of the word. […] Compare this extremely condensed account with the following quotation from Popper’s widely read ’Three Views of Human Knowledge’: ’Today the view of physical science founded by Oslander [^2] It refers of course to the german theologician ”Andreas Osiander”, who also wrote the preface to Copernicus’ De revolutionibus orbium coelestium; the typo is in the original article., Cardinal Bellarmino, and Bishop Berkeley [who according to Popper declare either all of physics, or particular physical theories to be instruments of prediction only] has won the battle without another shot being fired. Without any further debate about the philosophical issue, without producing any new argument the instrumentalistic view . . . has become the accepted dogma. It may well now be called the ’official view’. Considering this quotation we can at once accept the assertion that instrumentalism is a well-entrenched belief of many contemporary physicists and that it is accepted by them as a matter of course, and ’without debate’. Trying to explain this phenomenon we can point to the complete lack of historical perspective on the part of the younger generation who ’received quantum mechanics served up ready made’ ([87], 219) and who ’do not know the debates that settled [important] issues . . . for the fathers of the quantum theory’. But must we not say that the remarks of Popper and of some more recent critics of the Copenhagen Interpretation exhibit the very same ignorance or neglect of background knowledge? The realism-instrumentalism issue, after all, was one of the most central and most hotly debated topics of the older quantum theory and it was discussed to such an extent that its repercussions should have reached the ears even of a philosopher. Every paper of Bohr’s emphasizes that so far an instrument of prediction is all one can have and that this shortcoming is due to the absence of unrefuted hypotheses about the nature of atomic processes. How does this fit with Popper’s ’without any further debate’? - [1], page 279-280
@@ 3 Realism vs instrumentalism, in the interpretation of QM
The philosophical issue at the core of PKF’s reflections in the context of the foundations of QM, and in fact of much of his (early) production, is the debate between realism and instrumentalism. This ties directly with the issue of how physical theories should be understood and used, of their descriptive or representational role in addition to their use as tools for prediction of future experiences. We will not try to give a comprehensive summary of PKF’s position. We will only single out (and rather briefly) a couple of conclusions from his reflection, that we find useful for the foundations of QM and only discuss them in this context.
@@
Roughly, (scientific) realism is the view that the world and its constitutive objects, together with their properties, exist independently of our knowledge of them, and scientific theories somehow capture such independent reality. In this sense, scientific theories tell us ’what is real’ out there, whether we observe or know about it or not. Different versions of realism would then be identified in terms of how they articulate the correspondence between scientific theories and the world, and what they take this correspondence to imply about the theories. Establishing this correspondence requires relating the content of scientific theories to the results of our observations, our experiments, our experiences of the world, what can be labeled ’the phenomena’, but the correspondence will include also elements of the theories that have only indirect relation to phenomena. On the other hand, instrumentalism takes scientific theories to be pure instruments for prediction and organization of phenomena (i.e. again the result of direct observations), and judged only in terms of this, thus denying any relevance to elements of the theories which are not direct relation to such phenomena and, more generally, any meaningful sense in which scientific theories ’tell us what is real’ in the world. They are tools, they do not represent anything beyond what we experience  [^4] Of course, here we are simplifying, and somewhat ‘straw-manning’, both realism and instrumentalism as philosophical positions; both can be much more nuanced and complex, as well as less in contradiction with one another, and in general so they are, in the actual formulation and arguments of the scientists and philosophers who argue any version of such positions. This warning applies also to some of the considerations of PKF, and ours, in what follows..
@@
In the QM case as discussed by PKF, the debate between realism and instrumentalism is directly translated into the debate between the ’Copenhagen’ and the ’hidden variables’ interpretation of the QM formalism. This reflects the historical moment in which PKF was writing and the landscape of QM interpretations of the time. A certain standardized version of the QM interpretation coming from the Copenhagen school of Bohr, Heisenberg etc had become orthodoxy, and simplified from the richer variety and conceptual subtleties of the initial positions to become indeed a rather naive and dogmatic instrumentalist attitude. This reflected also a general discredit of conceptual and foundational work on physical theories that, together with increased specialization and separation of physics and philosophy, became dominant in the post-war period (for an historical analysis and some examples, see [13]). Similarly, realism in the context of QM is understood in a rather restricted sense. At the moment of PKF writings, the only well-developed alternative to the orthodox view was the deBroglie-Bohm mechanics [^5] For a survey of interpretations of QM, see [14], while for their individual tenets and history, see the references cited therein and the corresponding entire sin the Stanford Encyclopedia of Philosophy. based on position variables for particles understood as in classical Newtonian mechanics (but supplemented with the new ”pilot-wave” suggested by QM), and indeed the ”objective reality” threatened by QM and defended by ”realists” was the traditional one described by Newtonian mechanics. We will say more about modern perspectives on QM in the next section, but note here that there are more ways, nowadays, to be a realist in the context of QM (including both new proposed ontologies, based for example on ontic quantum states, like the many-worlds interpretation [15] and new ways of understanding realism itself [16, 17]), and that the Copenhagen perspective has evolved to be much richer and subtle than any straightforward instrumentalism (which can still be found defended with good arguments, though [18]).
@@
As said, there are a number of versions of scientific realism. PKF presents (some of) them in the introduction to [1], showing how the nature of scientific theories and their relation to ”facts” is in fact subtle, since theories are more complex than any naive picture of them as ”candidate portraits of reality” but also as ”systematization of observations” would want. This complexity should be properly appreciated in order to tackle the issue of realism vs instrumentalism in a fruitful manner, and this is indeed the preliminary goal of PKF’s analysis. In doing so, characteristically, PKF builds on the work of philosophers but even more on the practice of science and the insights coming from scientists. The remarks about the brilliance of Maxwell’s and Mach’s thinking, cited above, were written in this context. Again, what is particularly appreciated is the pragmatic, progress-oriented attitude of the scientists, opposed to the dogmatic, principles-first attitude of (some) philosophers, but also of scientists heavily influenced (in PKF’s view) by positivist philosophy.
@@
The first version of scientific realism therefore does not lead to a realistic interpretation for all theories, but only for those which have been chosen as a basis for research. It may be asserted (a) that the chosen theory has been shown to be true or (b) that it is possible to assume its truth, even though (ba) the theory has not been established or (bb) is in conflict with facts and established views. - [1], page 5
A second version of scientific realism assumes that scientific theories introduce new entities with new properties and new causal effects. […] a direct application of the second version of scientific realism (’theories always introduce new entities’) and a corresponding abstract criticism of’ positivistic’ tendencies are too crude to fit scientific practice. - [1], page 6-7
@@
The crudity of a purely philosophical approach becomes even clearer when we turn to a third version of scientific realism which is found in Maxwell, Helmholtz, Hertz, Boltzmann and Einstein. Naive realists - and many scientists and philosophers supporting the second version belong to this group - assume that there are certain objects in the world and that some theories have managed to represent them correctly. These theories speak about reality. The task of science is to discover laws and phenomena and to reduce them to those theories. […] Naive realism occurs in commonsense as well as in the sciences and it has been criticized in both. In the nineteenth century, the scientific criticism consisted in pointing out that theoretical entities and especially the theoretical entities of mathematical physics have a life of their own which may conceal the matter under examination. […] The accounts just given assume two different domains, or layers. On the one side we have phenomena, facts, things, qualities as well as concepts for the direct expression of their properties and relations. On the other side we have an abstract (quantitative) language in which the ’phantom pictures,’ i.e. scientific theories, are formulated. The pictures are correlated to the phenomena, facts, things, qualities of the first domain. Attention is paid to the language of the pictures or the ’theoretical language’, […] and one considers ways of modifying and improving it. Little attention is paid to the ’observation language’. - [1], page 8-11
@@
PKF is emphasizing that scientific theories have many components, with different roles and different and subtle relations to (empirical) facts, constructed via a mixture of a priori assumptions and hypotheses and of empirical feedback from observations, and a highly non-trivial intertwining of conceptual ingredients, mathematical language and observational inputs:
@@
A physical hypothesis does provide a guide and it also keeps the subject matter before our eyes. However, it makes us see the phenomena ’only through a medium’. Maxwell seems to fear that physical hypotheses may be imposed upon the phenomena without the possibility of checking them independently. As a result we cannot decide whether the phenomena are correctly represented by these hypotheses. Analogies avoid the drawbacks of mathematical formulae and of physical hypotheses. They are hypotheses in Mill’s sense of the word, i.e. assumptions about the nature of things which have been examined and have passed tests. They have heuristic potential, but they don’t blind us. […] The research instruments recommended by Maxwell differ from physical hypotheses not in content but in use: one follows the suggestions made by an analogy but checks them at every step by a comparison with independently described phenomena. Analogies are physical hypotheses restricted by such a process of checking and used with the thought of possible further restrictions firmly in mind. Their theoretical entities do not represent any real entities unless it turns out that the phenomena follow the hypothesis in every detail.- [1], page 11-12
@@
Moreover, it is a mistake, PKF argues, to believe that theories simply “represent” the world. This representational feature is questionable both for the directly observational components, which provide the basis for an instrumentalist interpretation, and for the more abstract components that we may take as indicating a ”hidden” (and maybe truer) physical reality, from a realist standpoint. Which aspects or parts of scientific theories do in fact represent is a choice grounded in physical understanding, not an immediately obvious one. One could say that it is concepts that represent, mediated, in case, by (mathematical) theories. This is a point that PKF makes also referring to Bohr’s position (see also [19]).
@@
Having appreciated such richness, one starts doubting about the key distinction on which the realism vs instrumentalism debate is built, i.e. the “double layer” constituted by ”theories” on the one hand and ”facts” on the other, as well as about the associated distinction between “theoretical language” and “observational language”, i.e. the “double language” model. This is the first target of PKF’s philosophical criticism. It becomes in particular a criticism of the instrumentalist position, which relies crucially on the above distinction in its insistence that only observations and empirical data should matter, in our development and then acceptance of physical theories, and that we do not have to seek anything more than that, in them. It is in fact part of PKF’s broader criticism toward empiricism and especially the positivist tradition (for more details beyond the above sketch, see again the introduction of part 1 in [1]).
@@
PKF argues that all concepts, including those used to formulate facts, are theoretical concepts. It is theories that influence what we see as ’facts’ and how we understand/interpret them. This is the well-known point of theory-ladenness of observations [20], by now accepted in philosophy of science (also thanks to PKF’s work, plus the earlier work of Popper and the later one of Laudan, to mention just a few). It follows that any sharp or absolute distinction between observations and theories, between empirical facts and theoretical constructions, is false and misleading. Thus, a pure instrumentalist position is either disingenuous or, in some sense, a realist position in disguise, in which one accepts some (elements of some) theories as pointing to something real and others as tools only. It is also, to some extent, an overly conservative position, because it ends up retaining, without questioning them enough, what are ultimately the conceptual theoretical elements of older scientific theories or philosophical traditions.
@@
The conclusion of PKF’s reasoning is that the observational language (which is the basis for empiricism, then instrumentalism, but also of certain realist positions, when one wants to retain as pointing out to something real, in our theories, those elements that more closely connect to the observational language) is also a theoretical one and should be treated as such.
@@
[…] we may now tentatively put forward our thesis I: the interpretation of an observation language is determined by the theories which we use to explain what we observe, and it changes as soon as those theories change. […] The consequences in whose light our thesis I should be judged, are (inter alia) the following: (1) According to thesis I, we must distinguish between appearances (i.e. phenomena) and the things appearing (the things referred to by the observational sentences in a certain interpretation). This distinction is characteristic of realism. (2) The distinction between observational terms and theoretical terms is a pragmatic (psychological) distinction which has nothing to do with the logical status of the two kinds of term. On the contrary, thesis I implies that the terms of a theory and the terms of an observational language used for the tests of that theory give rise to exactly the same logical (ontological) problems. […] (5) A theory (such as electrodynamics) may be understood even by a blind person. The only difference between a blind person and a seeing person consists in the fact that the first one uses a different part of the theory (or of the consequences of the theory) as his observation language. Hence, even a blind person may understand ’red’ and similar terms (of his theoretical language) and there is no reason why he should not be able to explain ’red’ to a seeing person ’by ostension’. This being so we cannot assume that when ceasing to be blind he automatically improves his knowledge of redness. It is to be admitted that he will now possess a new (and very effective) method of deciding (in the pragmatic sense of section 2) whether or not a given object is red. But just as the invention of a new microscope will change our notion of certain microscopic organisms only if it leads to new theories about them, in the same way the fact that our observer is now able to see red will lead him to a new notion of redness only if it leads him to new theories about red - and this need not be the case - [1], page 31-33
Not even everyday language should be taken for granted, accepted as indispensable element of our epistemology, and of our science, and taken as the basis for accepting a certain observational language.
It has been frequently asserted that the language in which we describe our surroundings, chairs, tables and also the ultimate results of experiment (pointer-reading) is fairly insensitive towards changes in the theoretical ’superstructure’. It seems somewhat doubtful whether even this modest thesis can be defended; first, because a uniform ’everyday language’ does not exist. The language used by the ’everyday man’ (whoever that may be) is a mixture of languages, as it were, i.e. it is a means of communication which has received its interpretation from various and often incompatible and obsolete theories. Secondly, it is not correct that this mixture does not undergo important changes: terms which at some time were regarded as observational elements of ’everyday language’ (such as the term ’devil’) are no longer regarded as such. Other terms, such as ’potential’, ’velocity’, etc., have been included in the observational part of everyday language, and many terms have assumed a new use. The fact that the pragmatic properties of some parts of the everyday language have remained unchanged may well be due to the fact that the people using these particular sections are not interested in science and do not know its results; after all, theories as such cannot influence linguistic habits. What can influence those habits is the adoption of theories by certain people. Yet it must be conceded that even the scientist who employs parts of everyday language for the purpose of giving an account of his experiments does not introduce a new use for familiar words such as ’pointer’, ’red’, ’moving’, etc. whenever he changes his theories. Does it follow, as is asserted in the stability thesis, that he is always talking about the same things and that he always employs the same interpretation of his observation language? Our foregoing analysis shows that this need not be the case. - [1], page 31
@@
This criticism of the double-layer and double-language model has immediate consequences for our understanding of QM, and for the debate between “Copenhagen” vs “hidden variables”.
@@
The first is the ensuing criticism of Copenhagen position, with its reliance on observations and its reduction of the possible reach of physical theories in general, supposedly learned from QM, to an account of them. The criticism is directed also at the conviction that our observations have to be expressed in classical terms, ie. those that are, upon scrutiny, inherited from classical Newtonian mechanics. That is, not only observations are theoretically loaded, as we agreed, but they are loaded of an older theory. There may be good reasons to accept this load, but they have to be argued for, not taken for granted. We note that this reliance on classical mechanics in providing the necessary observational language for QM is a key ingredient not only of the traditional Bohr/Copenhagen position, but also of its more modern incarnation [21]. According to PKF, this is a remnant of a hard-to-shake positivist influence. The danger of such uncritical acceptance is not only that it unnecessarily restricts our epistemological and scientific options, and dampens our creativity, but that it may quickly turn dogmatic. On the contrary, we should always leave open (and in fact actively pursue) the possibility of revising our theoretical structures, including those that constitute the basis of our observational language and those that follow from older theoretical schemes. PKF is very forceful in attacking this aspect of the Copenhagen orthodoxy, maybe insisting on a not-too-charitable interpretation of Bohr’s own words (which could be taken to be simply stressing the new lessons of QM and the need to take them seriously in future developments, not necessarily maintaining them unaltered):
@@
[…] the observation language of physics is a positivistic observation language whose interpretation is the same as the interpretation of classical physics before the advent of quantum mechanics. How can this be reconciled with the fact that classical physics is contradicted by the quantum of action? According to Bohr it can be reconciled by restricting the application of classical terms in a way which (a) ’provides room for new physical laws’ and especially for the quantum of action; which (b) still allows us to describe any possible experiment in classical terms; and which (c) leads to correct predictions. […] Bohr seems to assume that this will hold for any future theory of microscopic entities. Now it may be conceded that the laws of quantum mechanics do not admit of a straightforward interpretation on the basis of a classical model, as such a model would be incompatible either with the principle of superposition or with the individuality of the microscopic entities. It may also be conceded that as a matter of fact we do find it difficult (though by no means impossible) to form an intuitive picture of processes which are not dependent upon the classical framework. But from this psychological predicament we can by no means infer (assumption 1) that such intuitive understanding will never be possible. And it would be even less correct to assume on that basis that the concept of a non-classical process cannot be formed (assumption 2); for it is well known that we can form and handle concepts even of those things which we cannot readily visualize. […] according to Bohr it would even be a ’misconception to believe that the difficulties of the atomic theory may be evaded by eventually replacing the concepts of classical physics by new conceptual forms’, as there exist ’general limits of man’s capacity to create concepts’. How can this defeatist attitude be understood? I think that it can be understood if we explain more thoroughly the ideas upon which Bohr’s interpretation is based. The first idea is that the belief in classical physics has influenced not only our thinking but also our experimental procedures and even our ’forms of perception’. This idea gives a correct description of the effect which the continued use of a fairly general physical theory may have upon our practices and upon our perceptions: it will become increasingly difficult to imagine an alternative account of the facts. The second idea is inductivism. According to inductivism we invent only such theories as are suggested by our observations. Combined with the first idea inductivism implies that it is psychologically impossible to create non-classical concepts and to invent a non-classical ’conceptual scheme’. The third idea is the principle of pragmatic meaning. According to this idea, the use of classical methods and the existence of classical ’forms of perception’ imply that the observation language possesses a classical interpretation (see above). As a non-classical picture of the world would lead to an interpretation which is inconsistent with this classical interpretation, such a non-classical picture, apart from being psychologically impossible, would even involve a logical absurdity. […] As opposed to this it is sufficient to point out that even in a situation where all facts seem to suggest a theory which cannot any longer be maintained to be universally true, that even in such a situation the invention of new conceptual schemes need not be psychologically impossible so long as there exist abstract pictures of the world (metaphysical or otherwise) which may be turned into alternative interpretations. - [1], page 23-24

### v3

@@ 2 Relation between physics and philosophy, in the foundations of QM
[…] throughout the period of the older quantum theory Bohr underlined the ’formal’ character of the new ideas which arose in connection with the principle of correspondence: these ideas, we can read repeatedly in his earlier papers, lead to some surprising predictive successes, and they also establish some order in the steadily accumulating empirical material. However, they do not form a theory in the ordinary, or ’classical’ sense of the word, that is, they do not contain a coherent description of new and objective features. Nor can we say that they explain the relevant experimental facts. All this is emphasized by Bohr himself and not only once, but in almost every paper he writes. He also emphasizes that the dependence on classical ideas that is implied by the use of the correspondence principle characterizes ’the present state of science’, or ’the present state of our knowledge’, and not the essence of physics. It is even regarded as somewhat undesirable. In these early years Bohr’s aim seems to have been the gradual construction of an effective instrument of prediction which he hoped would act as a stimulus and as an aid towards the invention of a new atomic theory in the full sense of the word. […] - [1], page 279-280
@@ 3 Realism vs instrumentalism, in the interpretation of QM
The philosophical issue at the core of PKF’s reflections in the context of the foundations of QM, and in fact of much of his (early) production, is the debate between realism and instrumentalism. This ties directly with the issue of how physical theories should be understood and used, of their descriptive or representational role, in addition to their use as tools for prediction of future experiences. We will not try to give a comprehensive summary of PKF’s position. We will only briefly single out a couple of conclusions from his reflection, that we find useful for the foundations of QM and only discuss them in this context.
@@
Roughly, (scientific) realism is the view that the world and its constitutive objects, together with their properties, exist independently of our knowledge of them, and scientific theories somehow capture this independent reality. In this sense, scientific theories tell us ’what is real’ out there, whether we observe or know about it or not. Different versions of realism would then be identified in terms of how they articulate the correspondence between scientific theories and the world, and what they take this correspondence to imply about the theories. Establishing this correspondence requires relating the content of scientific theories to the results of our observations, our experiments, our experiences of the world, i.e. what can be labeled ‘the phenomena’. However, the correspondence will include also elements of the theories that have only indirect relation to phenomena. On the other hand, instrumentalism takes scientific theories to be pure instruments for prediction and organization of phenomena (i.e. again, the result of direct observations), and judged only in terms of this. Thus, it denies any relevance to elements of the theories which are not direct relation to such phenomena and, more generally, any meaningful sense in which scientific theories ‘tell us what is real’in the world. They are tools, they do not represent anything beyond what we experience  [^3] Of course, here we are simplifying, and somewhat ‘straw-manning’, both realism and instrumentalism as philosophical positions; both can be much more nuanced and complex, as well as more mutually conciliatory, in the actual formulation and arguments of the scientists and philosophers who argue for such positions. This warning applies also to some of the considerations of PKF, and ours, in what follows..
@@
In the QM case as discussed by PKF, the debate between realism and instrumentalism is directly translated into the debate between the ‘Copenhagen’and the ‘hidden variables’interpretation of the QM formalism. This reflects the historical moment in which PKF was writing and the landscape of QM interpretations of the time. A certain standardized version of the QM interpretation coming from the Copenhagen school of Bohr, Heisenberg etc had become orthodoxy, and simplified from the richer variety and conceptual subtleties of the initial positions to become indeed a rather naive, and dogmatic, instrumentalist attitude. This reflected also a general dismissal of conceptual and foundational work on physical theories that, together with increased specialization and separation of physics and philosophy, became dominant in the post-war period (for an historical analysis and some examples, see [23]). Similarly, realism in the context of QM is understood in a rather restricted sense. At the moment of PKF writings, the only well-developed alternative to the orthodox view was the deBroglie-Bohm mechanics [^4] For a survey of interpretations of QM, see [24], while for their individual tenets and history, see the references cited therein and the corresponding entries in the Stanford Encyclopedia of Philosophy. based on position variables for particles understood as in classical Newtonian mechanics (but supplemented with the new ”pilot-wave” suggested by QM), and indeed the ”objective reality” threatened by QM and defended by ”realists” was the traditional one described by Newtonian mechanics. We will say more about modern perspectives on QM in the next section, but note here that there are more ways, nowadays, to be a realist in the context of QM (including both new proposed ontologies, based for example on ontic quantum states, like the many-worlds interpretation [25] and new ways of understanding realism itself [26, 27]), and that the Copenhagen perspective has evolved to be much richer and subtle than any straightforward instrumentalism (which can still be found defended with good arguments, though [28]).
@@
As said, there are several versions of scientific realism. PKF presents (some of) them in the introduction to [1], showing how the nature of scientific theories and their relation to ”facts” is in fact subtle, since theories are more complex than any naive picture of them as ”candidate portraits of reality” but also as ”systematization of observations” would want. This complexity should be properly appreciated to tackle the issue of realism vs instrumentalism in a fruitful manner, and this is indeed the preliminary goal of PKF’s analysis. In doing so, characteristically, PKF builds on the work of philosophers but even more on the practice of science and the insights of scientists. His remarks about the brilliance of Maxwell’s and Mach’s thinking were written in this context. Again, what is particularly appreciated is the pragmatic, progress-oriented attitude of the scientists, opposed to the dogmatic, principles-first attitude of (some) philosophers, but also of scientists heavily influenced (in PKF’s view) by positivist philosophy.
@@
The first version of scientific realism therefore does not lead to a realistic interpretation for all theories, but only for those which have been chosen as a basis for research. It may be asserted (a) that the chosen theory has been shown to be true or (b) that it is possible to assume its truth, even though (ba) the theory has not been established or (bb) is in conflict with facts and established views.[…] A second version of scientific realism assumes that scientific theories introduce new entities with new properties and new causal effects. […] a direct application of the second version of scientific realism (’theories always introduce new entities’) and a corresponding abstract criticism of’ positivistic’ tendencies are too crude to fit scientific practice. - [1], page 5-7
@@
Naive realists - and many scientists and philosophers supporting the second version belong to this group - assume that there are certain objects in the world and that some theories have managed to represent them correctly. These theories speak about reality. The accounts just given assume two different domains, or layers. On the one side we have phenomena, facts, things, qualities as well as concepts for the direct expression of their properties and relations. On the other side we have an abstract (quantitative) language in which the ’phantom pictures,’ i.e. scientific theories, are formulated. The pictures are correlated to the phenomena, facts, things, qualities of the first domain. Attention is paid to the language of the pictures or the ’theoretical language’, […] and one considers ways of modifying and improving it. Little attention is paid to the ’observation language’. - [1], page 8-11
@@
PKF is emphasizing that scientific theories have many components, with different roles and different, subtle relations to (empirical) facts, constructed via a mixture of a priori assumptions and hypotheses and of empirical feedback from observations, and a highly non-trivial intertwining of conceptual ingredients, mathematical language and observational inputs.
@@
A physical hypothesis does provide a guide and it also keeps the subject matter before our eyes. However, it makes us see the phenomena ’only through a medium’. Maxwell seems to fear that physical hypotheses may be imposed upon the phenomena without the possibility of checking them independently. As a result we cannot decide whether the phenomena are correctly represented by these hypotheses. Analogies avoid the drawbacks of mathematical formulae and of physical hypotheses. They are hypotheses in Mill’s sense of the word, i.e. assumptions about the nature of things which have been examined and have passed tests. They have heuristic potential, but they don’t blind us. - [1], page 11-12
@@
Moreover, it is a mistake, PKF argues, to believe that theories simply “represent” the world. This representational feature is questionable both for the directly observational components, which provide the basis for an instrumentalist interpretation, and for the more abstract components that we may take as indicating a “hidden”(and maybe truer) physical reality, from a realist standpoint. Which aspects or parts of scientific theories do in fact represent is a choice grounded in physical understanding, not an immediately obvious one. One could say that it is concepts that represent, mediated, in case, by (mathematical) theories. This is a point that PKF makes also referring to Bohr’s position (see also [29]).
@@
Having appreciated such richness, one starts doubting the key distinction on which the realism vs instrumentalism debate is built, i.e. the “double layer” constituted by ”theories” on the one hand and ”facts” on the other, as well as about the associated distinction between “theoretical language” and “observational language”, i.e. the “double language”model. This is the first target of PKF’s philosophical criticism. It becomes, in particular, a criticism of the instrumentalist position, which relies crucially on the above distinction in its insistence that only observations and empirical data should matter in our development and acceptance of physical theories. This is a part of PKF’s broader criticism toward empiricism and especially the positivist tradition (for more details beyond the above sketch, see again the introduction of part 1 in [1]).
@@
PKF argues that all concepts, including those used to formulate facts, are theoretical concepts. It is theories that influence what we see as ’facts’ and how we understand/interpret them. This is the well-known point of theory-ladenness of observations [30], by now accepted in philosophy of science (thanks to PKF’s, Kuhn’s and Hanson’s work, plus the earlier work of Popper and the later work of Laudan, to mention just a few). It follows that any sharp or absolute distinction between observations and theories, between empirical facts and theoretical constructions, is false and misleading. Thus, a pure instrumentalist position is either disingenuous or, in some sense, a realist position in disguise, in which one accepts some (elements of some) theories as pointing to something real and others as tools only. It is also, to some extent, an overly conservative position, because it ends up retaining, without sufficient questioning, what are ultimately the conceptual theoretical elements of older scientific theories or philosophical traditions.
@@
The conclusion of PKF’s reasoning is that the observational language (which is the basis for empiricism, then instrumentalism, but also certain realist positions) is also a theoretical one and should be treated as such.
@@
[…] we may now tentatively put forward our thesis I: the interpretation of an observation language is determined by the theories which we use to explain what we observe, and it changes as soon as those theories change. […] (1) According to thesis I, we must distinguish between appearances (i.e. phenomena) and the things appearing (the things referred to by the observational sentences in a certain interpretation). This distinction is characteristic of realism. (2) The distinction between observational terms and theoretical terms is a pragmatic (psychological) distinction which has nothing to do with the logical status of the two kinds of term. On the contrary, thesis I implies that the terms of a theory and the terms of an observational language used for the tests of that theory give rise to exactly the same logical (ontological) problems. […] - [1], page 31-33
Not even everyday language should be taken for granted, accepted as indispensable element of our epistemology, and of our science, and taken as the basis for accepting a certain observational language. The reasons are that, first of all, there is no “uniform everyday language” but a mixture of languages “which has received its interpretation from various and often incompatible and obsolete theories”. Second, “terms which at some time were regarded as observational elements of ’everyday language’ (such as the term ’devil’) are no longer regarded as such. Other terms, such as ’potential’, ’velocity’, etc., have been included in the observational part of everyday language, and many terms have assumed a new use.”, i.e. what is part of everyday language changes with the adoption of new theories (quotes from [1], page 31).
@@
This criticism of the double-layer and double-language model has immediate consequences for our understanding of QM, and for the debate between “Copenhagen” and “hidden variables”.
@@
The criticism is directed in particular at the conviction (central in the Copenhagen position) that our observations have to be expressed in classical terms, i.e., those that are, upon scrutiny, inherited from classical Newtonian mechanics. That is, not only observations are theoretically loaded, as we agreed, but they are loaded by an older theory. There may be good reasons to accept this (implicit) theoretical basis, but they have to be argued for, not taken for granted. We note that this reliance on classical mechanics in providing the necessary observational language for QM is a key ingredient not only of the traditional Bohr/Copenhagen position, but also of its more modern incarnation [31]. According to PKF, this is a remnant of a hard-to-shake positivist influence. The danger of such uncritical acceptance is not only that it unnecessarily restricts our epistemological and scientific options, and dampens our creativity, but that it may quickly turn dogmatic. On the contrary, we should always leave open (and in fact actively pursue) the possibility of revising our theoretical structures, including those that constitute the basis of our observational language and those that follow from older theoretical schemes. PKF is very forceful in attacking this aspect of the Copenhagen orthodoxy, maybe insisting on a not-too-charitable interpretation of Bohr’s own words (which could be taken to be simply stressing the new lessons of QM and the need to take them seriously in future developments, not necessarily maintaining them unaltered):
@@
Bohr seems to assume that this will hold for any future theory of microscopic entities. Now it may be conceded that the laws of quantum mechanics do not admit of a straightforward interpretation on the basis of a classical model, as such a model would be incompatible either with the principle of superposition or with the individuality of the microscopic entities. It may also be conceded that as a matter of fact we do find it difficult (though by no means impossible) to form an intuitive picture of processes which are not dependent upon the classical framework. But from this psychological predicament we can by no means infer (assumption 1) that such intuitive understanding will never be possible. And it would be even less correct to assume on that basis that the concept of a non-classical process cannot be formed (assumption 2); for it is well known that we can form and handle concepts even of those things which we cannot readily visualize. […] according to Bohr it would even be a ’misconception to believe that the difficulties of the atomic theory may be evaded by eventually replacing the concepts of classical physics by new conceptual forms’, as there exist ’general limits of man’s capacity to create concepts’. - [1], page 23-24
The Copenhagen position, thus, according to PKF, gives a good example of how philosophical attitudes can both facilitate scientific progress and hamper it especially when they leave part of our epistemic tools (here, the observational language) unquestioned . Indeed, it follows directly from specific philosophical ideas:
The first idea is that the belief in classical physics has influenced not only our thinking but also our experimental procedures and even our ’forms of perception’. This idea gives a correct description of the effect which the continued use of a fairly general physical theory may have upon our practices and upon our perceptions: it will become increasingly difficult to imagine an alternative account of the facts. The second idea is inductivism. According to inductivism we invent only such theories as are suggested by our observations. Combined with the first idea inductivism implies that it is psychologically impossible to create non-classical concepts and to invent a non-classical ’conceptual scheme’. The third idea is the principle of pragmatic meaning. According to this idea, the use of classical methods and the existence of classical ’forms of perception’ imply that the observation language possesses a classical interpretation (see above). As a non-classical picture of the world would lead to an interpretation which is inconsistent with this classical interpretation, such a non-classical picture, apart from being psychologically impossible, would even involve a logical absurdity. […] As opposed to this it is sufficient to point out that even in a situation where all facts seem to suggest a theory which cannot any longer be maintained to be universally true, that even in such a situation the invention of new conceptual schemes need not be psychologically impossible so long as there exist abstract pictures of the world (metaphysical or otherwise) which may be turned into alternative interpretations. - [1], page 23-24